Circular No. 5920

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET KUSHIDA (1994a)

Ephemeris from the orbital elements on IAUC 5919:

1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 8 9^{h}45.^{m}38 + 3^{o}20.^{′}0 0.553 1.448 139.^{o}2 26.^{o}4 10.8
13 9 47.97 + 2 30.7 0.564 1.474 143.1 23.6 10.9
18 9 49.38 + 1 54.2 0.578 1.503 147.3 20.7 11.1
23 9 49.80 + 1 29.9 0.596 1.535 151.7 17.7 11.2
28 9 49.47 + 1 16.8 0.617 1.569 156.1 14.7 11.4
Feb. 2 9 48.59 + 1 13.7 0.642 1.606 160.5 11.8 11.6
7 9 47.38 + 1 18.9 0.672 1.644 164.6  9.2 11.8
12 9 46.07 + 1 30.8 0.707 1.684 167.7  7.2 12.0
17 9 44.86 + 1 47.5 0.746 1.726 168.9  6.3 12.2

Total visual magnitude estimate by R. Keen, Mt. Thorodin, CO (0.32m reflector): Jan. 11.31 UT, 10.7 (2′.5 coma).

NOVA CASSIOPEIAE 1993 R. W. Argyle and L. V. Morrison, Royal Greenwich Observatory, report

an accurate position for N Cas 1993 obtained by D. W. Evans with the Carlsberg Automatic Meridian Circle on La Palma: α = 23^{h}41^{m}47^{s}.228 ± 0^{s}.021, δ = +57^{o}31′00′′.76 ± 0′′.16 (equinox J2000.0, FK5 reference frame; epoch 1993.96). The corresponding 1950.0 position is α = 23^{h}39^{m}22^{s}.283, δ = +57^{o}14′22′′.55. The V magnitude was 6.12 (Dec. 16.75 UT). Only one observation was possible before the object passed into evening twilight.

Further photoelectric photometry: 1993 Dec. 30.75 UT, V = 7.08 ± 0.02 (H. Mikuž, Ljubljana, Slovenia; cf. IAUC 5914); 1994 Jan. 3.832, V = 7.16 ± 0.03, B–V = +0.43 ± 0.03, U–B = −0.64 ± 0.03 (D. Hanzl, N. Copernicus Observatory and Planetarium, Brno); Jan. 5.700, V = 7.19 ± 0.02, B–V = +0.34 ± 0.02 (Hanzl; comparison star SAO 35665, V = 7.24, B–V = +0.17).

Further visual magnitude estimates (cf. IAUC 5912): Jan. 6.05 UT, 7.1 (W. G. Dillon, Missouri City, TX); 6.11, 7.3 (C. E. Spratt, Victoria, BC); 8.88, 7.0 (S. Meister, Bulach, Switzerland); 9.17, 7.3 (A. Hale, near Orogrande, NM); 10.23, 7.1 (Spratt).

1994 January 11 Daniel W. E. Green